Circular No. 7800

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002J IN NGC 3464 D. Weisz and W. D. Li, University of California at Berkeley, report the

LOTOSS (cf. IAUC 7514) discovery of an apparent supernova on unfiltered CCD images taken with the KAIT on Jan. 21.4 (mag ∼ 16.8) and 22.4 UT (mag ∼ 16.6). SN 2002J is located at α = 10^{h}54^{m}38^{s}.36, δ = −21^{o}03′53′′.6 (equinox 2000.0), which is 23′′.5 west and 6′′.1 north of the nucleus of NGC 3464. KAIT images of the same field on Jan. 9.4 (limiting mag ∼ 19.0) and 15.4 (limiting mag ∼ 18.0) showed nothing at this position.

SUPERNOVA 2002C IN IC 3376 A. V. Filippenko and R. Chornock, University of California, Berkeley,

report that inspection of CCD spectra (range 330–1000 nm) obtained on Jan. 17 UT with the Keck-I telescope reveals that SN 2002C (IAUC 7793) is of type II, shortly after the explosion. The continuum is quite blue, and the Balmer lines (as well as He I 587.6-nm) have P-Cyg profiles.

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002C, obtained by P. Berlind on Jan. 19.49 UT with the Mt. Hopkins 1.5m telescope (+ FAST spectrograph), shows it to be a type-II supernova. P-Cyg lines of hydrogen are apparent on a blue continuum. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 7165 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 10 400 km/s.

SUPERNOVA 2002E IN NGC 4129 Matheson et al. also write that a spectrum of SN 2002E (cf. IAUC

7797), obtained by Berlind as above on Jan. 19.51 UT, shows it to be a type-II supernova past maximum: “Adopting the NED recession velocity of 1174 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 1300 km/s. The spectrum shows a relatively flat continuum (strongly contaminated by light from the host galaxy) with P-Cyg lines of hydrogen, but with weak absorption troughs. In addition, lines of [Ca II] (729.1, 732.4 nm) are present, indicating the beginning of the transition to the nebular phase. Other indicators of the nebular phase, such as the forbidden lines of oxygen, are not obviously apparent.”

2002 January 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green